### v2

# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

Abstract.
Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, verifiable computation, and secure finance. However, authoring ZK programs remains challenging: unlike conventional software development, ZK programming manifests a fundamental paradigm shift from imperative computation to declarative verification. This process requires rigorous reasoning about finite field arithmetic and complex constraint systems (which is rare in common imperative languages), making it knowledge-intensive and error-prone. While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and constraint-level reasoning, remains unexplored. To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities on ZK programming at three levels: language knowledge, algebraic primitive competence, and end-to-end program generation. Our evaluation of four state-of-the-art LLMs reveals that while models demonstrate strong proficiency in language syntax, they struggle when implementing and composing algebraic primitives to specify correct constraint systems, frequently producing incorrect programs. Based on these insights, we introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair. Experiments with GPT-o3 on Circom and Noir show substantial gains, with success rates improving from 20.29% to 87.85% and from 28.38% to 97.79%, respectively. With ZK-Eval and ZK-Coder, we establish a new basis for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance privacy computing.
copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
1. Introduction
Zero-knowledge proofs (ZKPs) are a powerful cryptographic primitive that allow one party to prove knowledge of a statement without revealing the underlying witness. In practice, this means proving that a computation or claim is correct while keeping sensitive inputs private. ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, verifiable computation, and emerging applications across finance and security domains (Sun et al., 2024; Gupta, 2025; Lavin et al., 2024; team, 2024). As demand for trustworthy computation grows, the ability to author, test, and verify ZKP programs has become a first-class concern in the software engineering community (Takahashi et al., 2025; Hochrainer et al., 2024; Xu et al., 2025; Pailoor et al., 2023; Jiang et al., 2025a).
Figure 1. Illustrative Comparison Between the Process of Mainstream and ZK Programming.
However, authoring ZK programs is challenging. Unlike mainstream programming, which builds on the von Neumann model and imperative execution, ZKP programs boil down to specifying and proving mathematical constraints over finite fields. In essence, this represents a paradigm shift from computation to verification. Its development requires reasoning about finite field arithmetic and constraint systems. As shown in Fig. 1 and Table 1, ZK programs do not describe how to execute a computation but how to prove its correctness by compiling constraints into arithmetic circuits. Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring algebraic primitives to specify constraints enforcing correctness. Even small omissions or miswirings can yield unsound or incomplete constraint systems, leading to errors that compilers rarely flag. This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025; Kolozyan et al., 2025; Chaliasos et al., 2025).
Table 1. Mainstream & ZK Program Comparison.
Aspect Mainstream ZK Program
Paradigm Imperative/OOP/Func Constraint Systems
Comp. Model von Neumann Arch Arithmetic Circuits
Purpose to Compute to Prove/Verify
Data Types Int/Float/Str/… Finite Field Elements
Mutability Dynamic Static
Control Flow If/Loop/Recursion Limited
Side Effects Yes No, Pure Math
Figure 2. Overview of ZK-Eval and ZK-Coder.
As a result, ZK programming demands significantly more time and expertise than mainstream software development, posing a steep barrier to entry for many practitioners. In contrast, recent advances in large language models (LLMs) have lowered this barrier in general-purpose languages by demonstrating strong code generation capabilities (Jiang et al., 2025b), achieving impressive results on benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021), and powering widely adopted tools like GitHub Copilot (GitHub, 2025) and Cursor (team, 2025). This contrast naturally raises the question: to what extent can modern LLMs assist in generating ZK code from natural language specifications? Answering this question requires more than surface-level benchmarking. Compared to programming in mainstream languages, success in ZK development hinges on three competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, (ii) mastery of algebraic primitives as atomic building blocks for constraint systems, and (iii) proficiency in modeling and specifying complex verification constraints. Existing benchmarks focus only on mainstream programming paradigms and overlook these requirements.
To measure this knowledge gap, this paper first introduces ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, primitive-level competence, and end-to-end program code generation (left panel of Fig. 2). We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why.
As will be detailed in § 3, applying ZK-Eval to four SoTA LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern. While models perform strongly on language knowledge, their accuracy drops significantly (below 30%) when reasoning about primitives or assembling verification constraints, highlighting a critical gap between syntactic knowledge and end-to-end generation. These results underscore both the potential of LLMs to democratize ZK development and the need for targeted augmentations to bridge this capability gap. Motivated by these findings, we present ZK-Coder (right panel of Fig. 2), which combines three key components: (i) a sketch layer (ZKSL) that explicitly models constraints using algebraic primitives, (ii) retrieval-augmented generation to ground specifications in verified implementations, and (iii) interactive refinement to fix constraint misspecifications. We observe that ZK-Coder achieves substantial gains, improving code generation success rates from 20.29% to 87.85% on Circom and from 28.38% to 97.79% on Noir (using GPT-o3). Ablation studies further confirm that sketching, guided retrieval and interactive refinement are indispensable; removing either component causes a significant performance drop (over 10%). In summary, our contributions are as follows:
  1. (1)
    We propose ZK-Eval, a three-stage evaluation pipeline for systematically assessing LLM capabilities in ZK code generation.
  2. (2)
    We introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair to improve reliability.
  3. (3)
    We evaluate the performance of four state-of-the-art LLMs on ZK-Eval and demonstrate how ZK-Coder substantially improves end-to-end code generation.
2. Background
2.1. Zero-Knowledge Proofs and Programming Frameworks
Zero-knowledge proofs (ZKPs) are cryptographic protocols that allow one party to prove to another that a statement is true without revealing any information beyond the validity of the statement itself. Formally, the following properties are desired for a correct ZKP program.
  1. (1)
    Completeness: If a statement is true, an honest prover can convince the verifier of its validity.
  2. (2)
    Soundness: If a statement is false, no dishonest prover can convince the verifier of its validity.
  3. (3)
    Zero-Knowledge: The verifier learns nothing beyond the validity of the statement itself.
These properties make ZKPs particularly useful in verifiable computation and privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024). However, ZK programming necessitates a fundamental paradigm shift from imperative computation to declarative verification. Rather than writing instructions to calculate a result, developers must model their high-level intent as an arithmetic representation $R(x,w)=1$ over a finite field, where $x$ denotes public inputs and $w$ denotes the private witness. This requirement to faithfully enforce logical intent through field-arithmetic constraints is notoriously arduous for humans; even minor errors in modeling can lead to under-constrained circuits (compromising soundness) or over-constrained systems (breaking completeness) (Pailoor et al., 2023; Wen et al., 2024a; Chaliasos et al., 2024; Kolozyan et al., 2025).
To mitigate this complexity, several programming frameworks have emerged, offering a diverse range of abstractions. We categorize these into two primary design paradigms:
  1. (1)
    Constraint-Oriented DSLs. Frameworks like Circom and Noir provide specialized syntax for expressing verification relations. Circom requires developers to think in terms of low-level circuit wiring and explicit signal management, which poses significant challenges for LLMs tasked with producing correct circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025). Conversely, Noir adopts a Rust-like syntax with strong typing, aligning more closely with the general-purpose coding corpora on which current LLMs are trained (Joel et al., 2024; Jiang et al., 2024a).
  2. (2)
    General-Purpose zkVMs. Emerging solutions like SP1 (Labs, 2025) and RISC0 (Contributors, 2025) allow developers to compile standard Rust code into a restricted instruction set, generating proofs over the execution trace within a virtual machine. While they abstract away the constraint system entirely, they often incur substantial performance overhead. We consider these an orthogonal approach to our work, as we focus on the direct generation of circuits.
In this work, we focus on Circom and Noir because they represent the most widely adopted frontends in both academia (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022; Pailoor et al., 2023; Wen et al., 2024a) and industry (e.g., in Ethereum rollups and Aztec). Furthermore, they occupy two complementary points on the abstraction spectrum: Circom tests an LLM’s ability to perform circuit-level specification and constraint wiring, while Noir evaluates how effectively models can adapt conventional coding proficiency to the verification-oriented paradigm. By studying both, we capture the core technical challenges of ZKP development while ensuring our findings have immediate practical relevance.
2.2. LLM-based Code Generation
Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024a, 2025b). These systems are reshaping engineering workflows, aiding in writing, testing, and documenting code by drawing on vast amounts of publicly available repositories and documentation (GitHub, 2025).
Framework. Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025a), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed. This process often involves multiple iterations, with the model refining its output based on user feedback or additional context. In this way, it can leverage off-the-shelf LLMs effectively without code-specific fine-tuning (Dong et al., 2025; Pan et al., 2025). However, to the best of our knowledge, there has been limited exploration of LLMs for code generation for low-resource programming languages (LRPL) or DSLs especially for ZKP DSLs.
Benchmark. Benchmarks are a crucial component of LLM-based code generation research, playing a vital role in understanding model performance and guiding improvements. Currently, their accuracy and reasoning ability are usually evaluated on standard benchmarks on mainstream programming languages such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021). However, evaluation tools for LRPLs and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025). Code generation for ZKP programs exemplifies the challenges and opportunities in this space. Though benchmarks for LRPLs (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
3. Measuring LLM Capability for ZK Code Generation
As discussed in § 1 and § 2.1, ZK programming follows a proof-oriented model: developers translate high-level intent into algebraic constraints over a finite field and produce proofs of satisfiability. This shift yields a unique programming workflow depicted in Fig. 1, where developers construct the verification relation by wiring together algebraic primitives (mathematical gates that enforce logical invariants over finite fields). While witness computation still resembles conventional imperative programming, the developer’s primary responsibility shifts to the rigorous specification of these verification constraints. To evaluate LLM proficiency in this critical domain, we assess two foundational capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ algebraic primitive competence.
  1. (1)
    Language and Toolchain Knowledge. Developers must master domain-specific syntax, strict typing rules, and the semantics of field operations to structure valid constraint specification and witness assignments. Weakness in any of these areas routinely leads to programs that cannot be compiled or produce an invalid relation that renders the proof meaningless.
  2. (2)
    Algebraic Primitive Competence. Algebraic primitives are the atomic building blocks of ZK constraint systems. Since arithmetic circuits lack native support for high-level operations, developers combine relational (e.g., comparators), logical (e.g., bitwise gates), and arithmetic (e.g., division, inversion) primitives to express logic. Competence in these atomic units is a prerequisite for constructing sound, higher-level application-specific constraint systems.
These capabilities constitute two core dimensions of end-to-end ZK program generation. Without language and toolchain knowledge, LLMs cannot produce compilable programs; without algebraic primitive competence, programs cannot faithfully encode intended relations. Since ZK programming requires integrating language-level fluency with deeper constraint modeling, existing mainstream benchmarks (Chen et al., 2021; Jain et al., 2024; Research, 2021) leave this integration unexplored. We introduce ZK-Eval to systematically measure both dimensions and their integration in end-to-end tasks.
3.1. ZK-Eval Benchmark
To realize this benchmark design, we structure ZK-Eval in three stages, each targeting one capability and their integration. We illustrate the design overview in Fig. 3.
Figure 3. The Three-Stage Evaluation Design of the ZK-Eval Benchmark.
3.1.1. Language and Toolchain Knowledge
We first evaluate whether LLMs possess the foundational knowledge required to author ZK programs. As shown in the left side of Fig. 3, sources are collected through web crawling of official documentation, including language specifications, tutorials, and API references (e.g., Circom’s circomlib (iden3, 2025) and the Noir’s stdlib). Additionally, we gather publicly available pedagogical examples from GitHub repositories (TEZCAN, 2025b; Sirin, ; Contributors, ; Thor, ) and online tutorials (TEZCAN, 2025a) to increase coverage of idiomatic usage. This yields 51 documents for Circom and 40 for Noir.
Based on this knowledge base, we then design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018). Inspired by earlier work on LLM-driven question generation (Liu et al., 2020; Abbasiantaeb et al., 2024; Long et al., 2024; Ji et al., 2025), our pipeline iteratively samples subsets of documents and provides them to a large model, which composes candidate questions and answers. This process is repeated until the full document collection has been exhaustively covered, producing a diverse set of drafts.
All generated questions are then subjected to rigorous expert validation. Three human reviewers independently assess each item for factual correctness, clarity, and alignment with the intended difficulty, with disagreements resolved through discussion to consensus. Cross-checking by the authors ensures consistency across the set. Topic coverage is also verified to ensure that all major aspects (as listed in the official documents) of the studied language are represented. We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library APIs), and compiler principles (translation from source code to arithmetic circuits). As shown in Fig. 3, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories. Representative examples are provided in the Supplementary Material and the full set of questions is provided in the Artifact.
3.1.2. Implementing and Using Algebraic Primitives to Specify Constraints
We assess the capability of LLMs to implement and utilize algebraic primitives as the building blocks of ZK constraint systems. Different from higher-level cryptographic (e.g., hash functions, signatures) and application-level (e.g., account-balance checks, order-matching validity) APIs available in external libraries, these encapsulate the core relational logic that developers must wire to define custom circuit behavior.
Technical Difficulty. We clarify that, constructing a benchmark that reflects the fundamental algebraic patterns essential to real-world ZK programs is challenging. In existing codebases, core relational logic is often entangled with application-specific boilerplate, making it difficult to isolate and evaluate the underlying algebraic reasoning. To decouple and isolate these foundational patterns, we systematically curate a set of primitives from widely adopted libraries. This process, illustrated in Fig. 3, draws from three complementary sources: circomlib (iden3, 2025), zk-kit (Nico Serrano, ), and Z3Py (Research, ).
  1. (1)
    Standardized Libraries. We first include canonical primitives from circomlib and zk-kit, which represent the production standard for routine circuit instantiation.
  2. (2)
    Agnostic Extraction. From these repositories, we manually extract and deduplicate algebraic primitives while filtering out application-specific composites. This ensures the benchmark captures the universal building blocks of ZK logic rather than application-specific patterns.
  3. (3)
    Aligned Completeness. To ensure comprehensiveness, we cross-reference our set with the Z3Py SMT-solver interface and incorporate primitives with similar structures, guaranteeing completeness for algebraic constraint specification.
The resulting benchmark consists of 35 algebraic primitives covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, such as addition, multiplication, integer division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum and multiplexers (see the full list in the Supplementary Material). This collection serves as a complete basis for constraint specification; as these algebraic primitives represent the atomic operations of arithmetic circuits, any non-trivial verification relation can be realized through their rigorous composition.
Practical Relevance. To check the practical relevance of our benchmark, we surveyed 37 production-grade ZK repositories comprising 408 source files. Our analysis confirms that the selected algebraic primitives are highly representative of real-world circuit logic: we achieve 97.30% and 94.59% coverage for core arithmetic and relational primitives, respectively, which are universal to production ZK systems. While logical (70.27%) and specialized composites (16.22%) appear less frequently, they remain mathematically essential for logical completeness and the construction of complex verification relations. Notably, it includes every one of the top 10 most frequently used primitives identified in the survey, establishing a representative foundation.
Task Setup. For each primitive, we author a concise task description that asks the LLM to implement the primitive’s functional logic and use it to specify the intended algebraic relation. Following established practice on test-case curation (Chen et al., 2021; Jain et al., 2024), we craft test-cases that validate both completeness (accepting valid witnesses) and soundness (rejecting invalid assignments). For finite primitives (e.g., logical xor), test-cases exhaustively enumerate the entire truth table. For primitives with larger domains, we combine automated random input generation with edge-cases (e.g., field overflows, zero-value or out-of-domain inputs), yielding at least 100 cases per primitive. To validate test rigor, we injected artificial errors into ground-truth implementations (like a mutation testing process); our tests detected 87.04% of these faults, confirming their effectiveness at catching subtle mistakes.
3.1.3. End-to-End ZK Program Generation
The final stage assesses LLMs to generate complete and correct ZK programs to verify computational statements. Building upon the first two stages, it evaluates the systemic integration of these skills to construct sound, multi-primitive constraint systems that faithfully encode high-level verification relations.
Benchmark Design Rationale. To ensure a rigorous and trustworthy assessment in the absence of a standardized ZK-specific generation dataset, we leverage canonical algorithmic baselines from the well-established HumanEval dataset. We posit that ZK programming represents a fundamental paradigm shift from imperative computation to algebraic verification rather than a shift in the application domain itself. Since the core logic of production-grade ZK applications (e.g., private finance, decentralized identity, and state-transition proofs) relies on universal algorithmic patterns (e.g., sorting, membership checks, and conditional branching), adapting these general algorithmic tasks allows us to isolate the gap on computation-to-verification translation without the confounding variables of application-specific requirements (e.g., application-specific API calling).
To reflect the unique semantics of ZK proofs, we systematically transform the original benchmark through three distinct adaptations, as illustrated in the right column of Fig. 3:
  1. (1)
    Compatibility Filtering. We exclude tasks involving string manipulation and floating-point arithmetic which are not supported by DSLs. Variable-length output tasks are also excluded since they’re incompatible with the static nature of arithmetic circuits. Their removal ensures that evaluation focuses on the model’s constraint-reasoning ability rather than DSL limitations. This yields a curated set of 68 standalone tasks suitable for ZK synthesis.
  2. (2)
    Verification Reformulation. We reformulate the natural language description of each computational task (defined as finding $S=P(I)$ for input $I$) into a verification problem. The LLM must generate a verification function $V(I,S)\in\{0,1\}$ that returns $1$ if and only if the candidate solution $S$ satisfies the intended relation for input $I$.
  3. (3)
    Oracle Adaptation. Correspondingly, the ground-truth oracle solutions are rewritten from computational functions ($S=P(I)$) to verification forms ($V(I,S)$). To ensure these serve as a reliable ground truth, we adopt a declarative implementation style; for instance, a sorting task is verified by checking permutation and monotonicity predicates rather than re-executing a sorting algorithm. To align with the semantics of ZKPs, these oracles are restricted to finite field arithmetic (mod $p$) and avoid non-deterministic execution paths. Each oracle was manually authored and cross-validated against the original computational benchmarks to ensure its reliability as a reference for test-case preparation.
Quality Assurance. Following established practices in benchmark construction (Yuan et al., 2023; Siddiq et al., 2023; Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022), we implement a rigorous validation process to ensure the Soundness and Completeness of generated programs. While traditional datasets provide only “Acceptance Cases” (Completeness), ZK programs require “Rejection Cases” to ensure the circuit is properly constrained (Soundness). Thus we implement test-case generators to produce input-solution pairs that a correct verifier must reject, ensuring soundness validation. We cross-validate task descriptions, adapted oracles, and test-case generators through manual review and injection of edge cases (e.g., all-zeros, field-overflows, out-of-domain inputs, etc.). Collectively, this process yields 100 test cases per task with an average distribution of 54.41% acceptance and 45.59% rejection cases. To quantify benchmark quality, we injected syntactic mutations into the ground-truth oracles (mutation-testing); our test suites detected 86.42% mutations, demonstrating high sensitivity to subtle errors.
3.2. Measurement Settings
Using ZK-Eval, we measure LLMs’ competencies on ZK code generation. Our study is guided by the following three research questions (RQs):
  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard library APIs of Circom and Noir?
  • RQ2: Algebraic Primitive Competence. Can LLMs reliably implement and use algebraic primitives, the building blocks of ZK programs, to specify correct constraints?
  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs on the intended computational relation?
Each RQ corresponds directly to one component of ZK-Eval. For RQ1, models are tested on curated multiple-choice questions covering language knowledge. For RQ2, they are instructed to implement and use algebraic primitives which are then automatically validated against acceptance and rejection test cases. For RQ3, they attempt end-to-end program generation to realize verification constraints for general computational problems, whose correctness is judged by soundness and completeness against dual test suites.
Model Selection. We evaluate 4 representative LLMs: OpenAI’s reasoning models GPT-o4-mini and GPT-o3, along with DeepSeek-V3 and Qwen3. This mix of reasoning and open-weight non-reasoning ones allows us to assess both family-specific and general trends. All models are run under identical prompts and budgets, with 10 samples per task to estimate Pass@1 rates and reduce variance.
Figure 4. Accuracy of LLMs and Human Experts on the MCQ Benchmark for ZK Language Knowledge.
Figure 5. LLM Performance on End-to-end ZK Benchmarks. Annotated with average accuracy.
Figure 6. Average Error Distribution of Primitive Implementations Across Languages, Types, and Causes.
3.3. (RQ1) Language & Toolchain Knowledge
Fig. 3.2 summarizes performance. Reasoning models perform best, with GPT-o4-mini reaching 88.1% and GPT-o3 87.2%. Open-weight non-reasoning models perform noticeably worse, with DeepSeek-V3 at 79.5% and Qwen3 at 78.9%. The human expert for reference scores 88.7%, similar to the reasoning models. Overall, LLMs demonstrate strong competence in ZK language knowledge, achieving accuracy within a few percentage points of the human expert.
Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-weight ones and perform similar to human experts.
We further examine performance across question categories and languages. LLMs achieve their highest accuracy on questions about basic syntax and API references, reflecting familiarity with fundamental constructs and widely used library templates. In contrast, accuracy drops on advanced topics such as traits, signal taggings, and generics, which extend expressiveness but are less commonly encountered. The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits, indicating that they struggle with deeper aspects of arithmetic circuits and constraint semantics on finite fields. Notably, there is no significant difference between Circom and Noir, suggesting that model competence reflects general exposure to ZK languages rather than language-specific familiarity.
Finding ②: LLMs excel at surface-level knowledge such as syntax and API references, but exhibit weaknesses in advanced language features and compiler principles.
These findings are consistent with the nature of the knowledge base: official documents, public tutorials, and examples are all widely accessible training material to LLMs. Thus, it is anticipated that LLMs can achieve strong performance on such content. This suggests that (i) explicit augmentations on language grammars or official documentations may be unnecessary (ii) LLMs exhibit strong potential for ZK programming tasks but require further evaluations to reveal the bottlenecks.
3.4. (RQ2) Algebraic Primitive Competence
Language knowledge enables syntactic fluency, but ZK programming fundamentally requires the ability to specify algebraic constraints. We next assess whether LLMs can use or implement algebraic primitives, the atomic elements of constraint systems. Specifically, model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes. From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) functional errors, where code compiled but failed functional correctness tests.
Fig. 6 shows the results and performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir. Comparison scored lower (37% and 20%), while Arithmetic and Composites dropped to extremely low accuracy. This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into specifying constraints with algebraic primitives, LLMs largely fail to produce testable, correct implementations. This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by the inherent programming paradigm shift (Sheybani et al., 2025).
Finding ③: On Circom and Noir languages, LLMs perform consistently weak on implementing and using algebraic primitives to specify constraints across all categories.
Typing & resolution errors are the most frequent. Noir enforces strict type matching, disallowing implicit casts between fields and integers, while Circom distinguishes between variables and signals. Models frequently mix types, reference undeclared identifiers, or redefine variables, causing compilation failures. Signal handling errors, unique to Circom, are the next most common, including declaring signals inside control-flow, reassigning them multiple times, or using them before initialization. Constraint enforcement errors also recur: models generate invalid constraint forms such as non-quadratic constraints (e.g., x * y * z === 1) or use off-circuit assertions. These issues stem from ZK-specific semantics absent in conventional languages.
Performance further varies by primitive type, reflecting library support and algebraic complexity. Logical operations and comparisons, being boolean primitives or supported by libraries, are most tractable. Arithmetic and Composites, such as exponentiation, modulo, floor division, or composites like sum and product, lack native support and must be implemented from scratch. These tasks demand reasoning over field arithmetic, careful handling of inverses and negatives, and explicit soundness constraints. The very low accuracy here underscores the inability of current models to generate implementations beyond those directly provided in libraries.
Finding ④: LLMs handle boolean and comparison primitives better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
3.5. (RQ3) End-to-End Generation
Fig. 5 summarizes the results on the end-to-end generation benchmark. Compared to language knowledge and primitive-level tasks, performance drops substantially across all models. For Circom, accuracy is very low: GPT-o3 reaches only 20% semantically correct and 22% syntactically valid programs, while Qwen3 drops to 5% and 9%. Noir results are somewhat higher but still limited, with GPT-o4-mini at 32% and 40% and Qwen3 at 10% and 19%.
Several trends emerge from these results. First, Noir tasks consistently yield higher success rates than Circom, likely due to Noir’s simpler type system and cleaner syntax, which reduce opportunities for errors such as signal misuse or type mismatches that are pervasive in Circom. Second, reasoning models substantially outperform open-weight ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek-V3 and Qwen3 across both languages, confirming the model capability difference already observed in earlier tasks. Finally, the gap between syntactic and semantic correctness underscores that being compilable alone is insufficient for assessing ZK program quality: many outputs are compilable but suffer from under- or over-constrained relations.
Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-weight ones; Noir is easier than Circom; many programs compile but remain incorrect.
This reveals the fundamental challenge of ZK programming: it requires not only language fluency and primitive competence, but also the ability to architect multi-primitive constraint systems that faithfully encode complex verification relations — a skill that current LLMs have not yet mastered.
4. Augmenting LLMs for ZK Code Generation
Based on the findings and gaps revealed from our empirical study in § 3, we introduce ZK-Coder, an agentic framework to augment end-to-end ZK code generation. We first summarize the design principles derived from our findings, then present the system.
4.1. Design Principles Learned from the Findings
Findings ① and ② show that LLMs possess strong knowledge of ZK languages’ syntax, semantics, and standard library APIs, achieving near expert-level accuracy. Overall, we interpret that knowledge has been internalized during contemporary LLM’s pretraining stage, eliminating the need for grammar prompting or documentation access in our augmentation. However, important challenges remain that motivate our further exploration.
In Findings ③ and ④, we observe that this knowledge does not translate into reliable primitive-level generation. Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement. Finding ⑤ also reveals these failures in end-to-end generation. To bridge this gap, ZK-Coder introduces a sketch layer that decouples abstract constraint modeling from low-level DSL implementation. By first formalizing relations in this intermediate representation, the framework can precisely map identified primitives to verified circuit templates and usage hints via retrieval-augmented generation (RAG) (Gao et al., 2023). Loosely speaking, this trades brittle memorization for grounding. Once the logical “what” is defined in the sketch, retrieval provides the exact “how” for DSL-specific wiring, collapsing the search space and enabling the reliable composition of even rare algebraic snippets.
Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to constraint-misspecification or under/over-constrained logic. These failures directly undermine the reliability and necessitate mechanisms to improve validity and faithfulness. To mitigate these failures, ZK-Coder introduces an interactive refinement loop. The LLM proposes a program, queries the compiler as an oracle to check constraint specification validity, and, once passes, self-supplies acceptance/rejection test cases to validate relation faithfulness. By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until these checks pass. This interaction transforms failures into actionable feedback and guides candidates toward valid, faithful programs.
4.2. ZK-Coder
Based on these design principles, we developed ZK-Coder, illustrated in Fig. 7. ZK-Coder augments LLMs with a structured pipeline that bridges natural-language problem descriptions and executable ZK programs. It consists of three main stages: ➀ constraint formulation, ➁ sketch-guided analysis and retrieval, and ➂ interactive generation and repair.
Figure 7. Illustrative Overview on the Design of ZK-Coder.
4.2.1. Constraint Formulation in Sketch Language.
$\begin{array}[]{rrcl}\textsc{Program}&P&::=&\texttt{def}\ \texttt{verify}(\texttt{engine},\ x_{1}\!:\!\tau_{1},\ \ldots,\ x_{n}\!:\!\tau_{n}):\ S^{*}\\[2.0pt] \textsc{Type}&\tau&::=&\texttt{Field}\mid\texttt{Bool}\mid\texttt{Vec[}\tau\texttt{]}\\[2.0pt] \textsc{Stmt}&S&::=&x=E\mid\texttt{engine.add}(\Phi)\mid\texttt{for}\ x\ \texttt{in}\ R:\ S^{*}\mid\texttt{if}\ \Phi:\ S^{*}[\texttt{else}:\ S^{*}]\mid\texttt{return engine}\\[2.0pt] \textsc{Range}&R&::=&\texttt{range}(E)\mid\texttt{range}(E,E)\\[2.0pt] \textsc{Formula}&\Phi&::=&A_{0}\mid\texttt{not}\ \Phi\mid\Phi\ \texttt{and}\ \Phi\mid\Phi\ \texttt{or}\ \Phi\mid\texttt{all}(\Phi\ \texttt{for}\ x\ \texttt{in}\ R)\mid\texttt{any}(\Phi\ \texttt{for}\ x\ \texttt{in}\ R)\\[2.0pt] \textsc{Atom}&A_{0}&::=&E\ \mathrel{\bowtie}\ E\mid G(E_{1},\ldots,E_{n})\\[2.0pt] \textsc{Expr}&E&::=&n\in\mathbb{Z}\mid\texttt{True}\mid\texttt{False}\mid x\mid(E)\mid E\ +\ E\mid E\ -\ E\mid E\ *\ E\mid E\ /\ E\mid E\%E\mid E//E\mid\ldots\\[2.0pt] \textsc{RelOp}&\bowtie&::=&\texttt{==}\mid\texttt{!=}\mid\texttt{<}\mid\texttt{<=}\mid\texttt{>}\mid\texttt{>=}\mid\texttt{and}\mid\texttt{or}\mid\texttt{xor}\\[2.0pt] \textsc{Primitives}&G&::=&\texttt{distinct}(E)\mid\texttt{select}(E,E,E)\mid\texttt{abs}(E)\mid\texttt{sum}(E\ \texttt{for}\ x\ \texttt{in}\ R)\mid\ldots\end{array}$
Figure 8. The Abstract Grammar of the ZK Sketch Language.
The first stage of ZK-Coder translates natural-language requirements into a structured specification of verification constraints. Drawing on the success of planning-based code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024b; Wen et al., 2024b), we introduce the ZK Sketch Language (ZKSL), a Python-embedded Domain Specific Language (eDSL) designed to bridge the gap between high-level intent and low-level circuit implementation.
ZKSL is architected to allow LLMs to formalize constraint logic using familiar Pythonic syntax. This isolates the mathematical “what” from the DSL-specific “how” by requiring the explicit modeling of constraints over typed inputs to mirror the verification nature of ZK programs. As formally defined [^1] We also provide concrete running examples in the Supplementary Material. in Fig. 8, a ZKSL Program is structured around a verify function that accepts these typed inputs and executes a sequence of constraint-declaration statements via engine.add($\Phi$). These formulas are composed of relational atoms that correspond to the algebraic primitives identified in § 3.1.2. Crucially, all operations within ZKSL are defined over finite fields (modulo $p$) to faithfully reflect the mathematical foundation of ZK systems. This enables LLMs to express complex verification relations using standardized constructs, serving as a direct logical precursor to the final target code. It allows LLMs to decouple algebraic reasoning from the implementation-level details (e.g., signal wiring and constraint encoding) of specific target DSLs.
ZK-Coder prompts the LLM to generate these constraint specification sketches directly from problem descriptions. To ensure the integrity of the sketch, the framework performs automated syntactic and type/arity consistency checks. If a sketch is found to be invalid, the error trace is fed back to the model for iterative refinement. By serving as a logical representation, ZKSL decouples the complex task of algebraic constraint reasoning from the low-level specifics of ZK DSLs (e.g., signal wiring and constraint specification formats), reducing the overall generation complexity.
4.2.2. Sketch-guided Analysis and Retrieval.
RAG is a powerful technique for grounding LLM generation in low-resource or domain-specific contexts (Joel et al., 2024). ZK-Coder utilizes the ZKSL sketch as a structural query to bridge the gap between abstract constraint specifications and code implementations. Since a ZK program is defined by a set of verification constraints composed of atomic algebraic primitives, the framework traverses the sketch to extract each relation declared via engine.add($\cdot$). To ensure precise retrieval, operators are canonicalized by normalizing aliases (e.g., >= becomes GreaterEqThan) and distinguishing variants based on their arity and field-typing (e.g., differentiating XOr for Booleans vs. Fields). This stage isolates the constituent units within each constraint to determine the precise implementation hints required by the target DSL.
To support this stage, we curate a knowledge base of verified implementations and idiomatic usage patterns for the algebraic primitives defined in § 3.1.2. Instead of text-similarity ranking, ZK-Coder then performs exact-match retrieval for each identified primitive. These patterns are emitted in the original ZKSL specification order to preserve the program’s logical flow during final code generation. This structured retrieval effectively grounds the generation process, collapsing the search space and significantly curbing the model’s tendency to hallucinate incorrect constraint semantics or signal-handling logic.
4.2.3. Interactive Refinement Loop.
While the sketch provides a logical blueprint, our findings in § 3.4 and § 3.5 reveal a persistent gap: many generated programs successfully satisfy DSL grammar rules but fail to faithfully encode the intended verification relation. Unlike mainstream software where errors often manifest as runtime crashes, ZK programs frequently yield under-constrained circuits or over-constrained systems, which silently accept invalid witnesses or sometimes reject valid ones. To bridge this gap, we implement an Interactive Refinement Loop as follows.
  1. (1)
    Constraint Construction Refinement. The candidate program is first submitted to the target compiler (e.g., Circom or Noir) to address DSL-specific failures identified in our study, such as signal mismanagement, illegal signal declarations in control flow, or constraint-encoding violations. We extract these diagnostic traces to iteratively refine the constraint wirings until they’re valid or reach the retry iteration threshold $N_{1}$ ($N_{1}$ is set as 8 in our experiments).
  2. (2)
    Relation Faithfulness Refinement. Once compilable, the framework validates constraint faithfulness. The agent generates representative “Acceptance” and “Rejection” test cases to verify that the circuit correctly enforces the intended relation, while evaluation test cases remain hidden to prevent data-leakage. If any test case fails, the counterexample is fed back to re-examine the algebraic constraints. This forms another refinement loop, and we set the retry threshold $N_{2}$ as 3 in our experiments.
The above processes repeat until the program passes all functional checks or the budgets are exhausted, practically converting failures into guided refinements without manual intervention.
5. Evaluation
We evaluate ZK-Coder on the task of generating zero-knowledge programs from natural language specifications. Our evaluation is structured around the following research questions, with a focus on end-to-end generation. To clarify, we do not re-evaluate language knowledge (RQ1) or primitive-level coding (RQ2): the former is established in § 3.3, and the latter is handled by our sketching and retrieval mechanism in § 4. Nevertheless, we include an ablation study (RQ5) to quantify the contribution of each component.
  • RQ4 (End-to-End Performance): How does ZK-Coder perform compared to baseline methods in generating correct ZK programs?
  • RQ5 (Ablation Study): What is the effect of each design component, including the ZK sketch language, constraint-guided retrieval, and the interactive repair loop?
  • RQ6 (Robust Generalization): How well does ZK-Coder generalize to novel, complex computational tasks and real-world production-grade ZKP implementations?
  • RQ7 (Failure Analysis): When ZK-Coder fails, what are common failure patterns, and how might they inform future research?
5.1. Experimental Setup
Our experiments are organized according to the research questions outlined above. For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the same four models from our measurement study. This comparison establishes how well ZK-Coder outperforms baseline methods in generating correct and provable ZK programs.
For RQ5 (Ablation Study), we conduct an ablation study on all four models, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop. This aims to quantify their individual impact on overall performance.
For RQ6 (Robust Generalization), we construct a new dataset following the adaptation methodology of ZK-Eval using LiveCodeBench. Unlike the simple and potentially contaminated HumanEval, this benchmark contains harder, contamination-free, and more recent tasks. We evaluate whether ZK-Coder continues to maintain high performance under more challenging scenarios.
Finally, for RQ7 (Failure Analysis), we analyze cases where ZK-Coder fails, categorizing the dominant error patterns and their underlying causes. This analysis highlights limitations of the current pipeline and suggests avenues for future research.
All experiments are conducted under a fixed repair budget (see § 4.2.3). The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
5.2. RQ4: End-to-End Performance
We evaluate ZK-Coder’s end-to-end performance. To the best of our knowledge, since no prior work directly addresses ZK program generation from natural language, we establish a simple but meaningful baseline: prompting the model to generate ZK code directly from task descriptions, with a summary of the target DSL’s grammar and few-shot examples. This setting captures the raw capability of LLMs as a lower bound. We provide more baselines in § 5.3.
We meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial constraint sketch (in ZKSL) produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases. The final overall accuracy aggregates across the full pipeline, reflecting end-to-end program generation accuracy. In addition, we report the average token cost for each model, which quantifies the spending.
Table 2. Comparative Per-stage Performance (Accuracy, Tokens) of ZK-Coder and Baseline.
Language Model ZK-Coder Per-stage Performance ZK-Coder Overall Baseline
(1) Sketch (2) Repair (3) Program Accuracy Avg. Accuracy Avg.
Correctness Pass Rate Correctness (Pass@1) Token (Pass@1) Token
Circom GPT-o4-mini 94.12% 91.72% 96.25% 83.38% 2350.31 17.35% 436.98
GPT-o3 97.30% 94.41% 95.64% 87.85% 2155.39 20.29% 433.35
Deepseek V3 72.15% 92.77% 74.10% 49.60% 2958.72 7.35% 325.28
Qwen 3 68.12% 79.07% 78.78% 42.43% 4979.72 4.90% 162.83
Noir GPT-o4-mini 94.12% 97.97% 96.38% 88.87% 648.95 32.21% 152.18
GPT-o3 97.94% 99.85% 100.00% 97.79% 750.63 28.38% 191.72
Deepseek V3 74.88% 93.36% 94.31% 68.66% 628.69 16.18% 100.91
Qwen 3 66.84% 91.88% 89.14% 54.74% 1260.43 11.27% 104.24
Table 2 summarizes the comparative performance of ZK-Coder. Across all tested models and languages, ZK-Coder achieves substantial gains over the baseline. On Circom, GPT-o3 leads with 87.85% accuracy, while open-weight models (DeepSeek V3, Qwen 3) improve from below 10% baseline to 49.60% and 42.43%, respectively. On Noir, success rates are higher overall, likely due to its Rust-like syntax being better aligned with pretraining data. GPT-o3 still leads with 97.79% accuracy, and GPT-o4-mini achieves 88.87%. Open-weight models also show strong gains: Qwen 3 improves from 11.27% to 54.74%, and DeepSeek V3 from 16.18% to 68.66%.
The per-stage breakdown highlights the robustness of the pipeline. OpenAI models show exceptional “zero-shot” capabilities in sketching constraints, while open-weight models exhibit lower correctness rates, suggesting that the leap from natural language to abstract constraint logic remains a challenge for these models. The repair loop maintains strong success rates, exceeding 90% in most settings, though Qwen 3 on Circom achieves 79.07%, indicating some difficulty in recovering from initial errors. Final program correctness is high across most configurations, though DeepSeek V3 and Qwen 3 show comparatively lower rates on Circom, suggesting challenges in implementing correct constraint logic. Overall, the integration of sketching, repair, and semantic validation yields consistently executable and provable programs across the majority of settings.
We also note that ZK-Coder incurs a higher token cost compared to the baseline, which is expected since our agentic pipeline includes sketching, RAG and interactive repair. However, the cost remains modest in practice, with each task requiring fewer than 5,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
Answer to RQ4: ZK-Coder delivers a significant improvement in end-to-end performance over the baseline method. These results confirm that our pipeline substantially enhances the performance on LLM-generating ZK programs.
5.3. RQ5: Ablation Study
Figure 9. Accuracy Comparison of ZK-Coder Ablation Variants (Averaged Across Models).
To assess the role of each major component in ZK-Coder, we conduct an ablation study where individual modules are disabled or replaced. The following variants are considered:
  • No RAG (V1): Disables sketch-guided RAG for algebraic primitives, forcing the model to rely solely on its internal knowledge.
  • No Sketch (V2): Removes the intermediate ZK constraint sketching and directly prompts the model to output full programs with RAG and the interactive repair loop.
  • No Compile-Repair (V3): Eliminates the interactive repair loop for compiler-identified errors.
  • No Execute-Repair (V4): Retains compiler-guided repair but disables execution and repair.
  • Similarity-based RAG (V5): Replaces our sketch-guided retrieval with a naive text-embedding similarity retrieval on the task description.
  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on test-and-repair.
Figure 9 presents the Pass@1 accuracy for each ablation variant across Circom and Noir. The results show that every component plays a critical role in the overall success of ZK-Coder. Starting from a full-system average of 65.66% on Circom and 76.87% on Noir, we observe that removing RAG on primitives (V1) causes a significant drop to 53.04% and 72.00% respectively. Disabling the constraint sketching stage (V2) causes further degradation (46.37% on Circom, 64.63% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint modeling. The most severe drop arises from removing compile-repair (V3), with accuracy falling to 26.04% on Circom and 27.12% on Noir, demonstrating that interactive repair is indispensable. Disabling execute-repair (V4) also leads to measurable reductions (64.48% on Circom, 76.00% on Noir), confirming that semantic validation contributes to robustness. Substituting primitive-in-sketch-guided retrieval with a text-similarity-based strategy (V5) yields weaker performance (51.42% on Circom, 74.59% on Noir), suggesting that sketch-guided retrieval is more effective particularly for Circom. Finally, “Only Repair” (V6) achieves only 43.31% on Circom and 59.80% on Noir, showing that iterative refinement alone cannot overcome the “computation-to-verification” gap. Compared with V6, the full system achieves an up to 22.35% performance leap with modest token cost (20.46%-29.39%), confirming that constraint sketching and template retrieval outperform blind repair cycles.
Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, sketch-guided RAG, and interactive repair. Removing or weakening any of these components causes noticeable accuracy drops.
5.4. RQ6: Robust Generalization
Our adapted HumanEval serves as a controlled benchmark to evaluate the core computation-to-verification shift. To evaluate the robustness of ZK-Coder across varying levels of difficulty and practical contexts, we extend our analysis to two additional datasets: (i) LiveCodeBench (Jain et al., 2024), which focuses on high-complexity algorithmic challenges to assess the limits of the framework’s reasoning; and (ii) a set of Production Case Studies curated from 37 open-source ZKP repositories. To ensure the case studies are representative and avoid selection bias, we employed a reproducible filtering process targeting high-frequency circuit patterns such as Merkle verification, state nullification, etc. Details are provided in the Supplementary Material.
Table 3. Generality Study on ZK-Coder’s Code Generation Performance (Averaged Across Models).
Language LiveCodeBench Production Coding Pattern Case Studies
ZK-Coder Repairing Baseline Simple Baseline ZK-Coder Repairing Baseline Simple Baseline
Circom 44.10% 14.07% 1.72% 90.83% 83.33% 27.50%
Noir 56.93% 33.95% 9.89% 92.09% 89.17% 31.67%
We establish two baselines: the Simple Baseline (direct generation) and the Repairing Baseline (adding the interactive repair loop). Table 3 presents the results. On LiveCodeBench, ZK-Coder achieves a dramatic performance leap, reaching 44.10% on Circom and 56.93% on Noir, significantly outperforming the Repairing Baseline (14.07% and 33.95%). On Production Case Studies, ZK-Coder reaches near-perfect success rates (90.83% on Circom and 92.09% on Noir). While the Repairing Baseline also performs strongly, this high baseline is expected, as these studies focus on standardized, commonly used verification patterns in a controlled context where the high-level logic is well-defined. Nevertheless, ZK-Coder achieves a noticeable gain by bridging the “last mile” of correctness, reducing subtle field-level constraint errors that persist in the baselines.
Answer to RQ6: Results from LiveCodeBench and production case studies confirm that ZK-Coder generalizes effectively across a spectrum of challenges, from high-complexity verification logic to standardized production coding patterns.
5.5. RQ7: Failure Analysis
Table 4. Distribution of ZK-Coder’s Code Generation Failures (Averaged Across Models).
Language ➀ Repair ➁ Sketched ➂ False Acceptance ➃ False Rejection ➄ Mixed False
Budget Exceed Constraint Incorrect (Under-constrained) (Over-constrained) Acc / Rej
Circom 31.67% 41.56% 1.08% 19.78% 5.92%
Noir 12.60% 73.07% 1.67% 10.89% 1.78%
Table 4 reports the distribution of ZK-Coder’s failure cases. Each column corresponds to a distinct failure mode. Repair Budget Exceed indicates cases where the repair budget was exhausted without producing any compilable program. Sketch Incorrect denotes that the generated constraint sketch failed hidden test cases, suggesting an incorrect specification of the verification constraints. False Acceptance corresponds to programs that erroneously accept invalid witnesses, while False Rejection refers to programs that incorrectly reject valid witnesses. Mixed False Acc / Rej captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
The data reveal that the two dominant sources of failures are ➀ and ➁. In both Circom and Noir, these two together account for over 70% of failures. Failure ➀ often arises when the model is unable to generate any compilable program, which is due to gaps in pretraining exposure to ZK-specific constructs. Failure ➁, on the other hand, reflects errors in correctly modeling the verification constraints from natural language descriptions. The less frequent categories (➂, ➃, ➄) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit. These errors indicate local implementation mistakes rather than systemic failures of the pipeline.
Answer to RQ7: In summary, Repair Budget Exceed and Sketched Constraint Incorrect constitute the most critical bottlenecks. The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise verification constraints.
6. Related Work
LLM for DSL Code Generation.
Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024; Giagnorio et al., 2025). To mitigate these difficulties, grammar prompting (Wang et al., 2023), RAG techniques (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024), iterative refinement (Ding et al., 2024), and knowledge transfer methods (Cassano et al., 2024; Mora et al., 2024) have been proposed. While most of these methods (except knowledge transfer methods, which target mainstream computational languages) can be adapted for ZKP code generation, our ablation study (§ 5.3) shows they remain ineffective. ZK-Coder advances this area by designing a system tailored to the constraint paradigm of ZKPs.
Code Generation Benchmarks.
Most benchmarks for LLM-based code generation focus on general-purpose languages (Jain et al., 2024; Chen et al., 2021), with limited support for DSLs. A recent survey (Joel et al., 2024) reviews existing DSL benchmarks and emphasizes the lack of standardized evaluation resources, highlighting the need for benchmarks that capture DSL-specific paradigms, idioms, and practical applications. In the case of ZKPs, to our knowledge, no systematic benchmark exists for ZKP code generation that reflects ZK-specific requirements, namely proving a computation, a static arithmetic-circuit model, and correctness defined by both acceptance and rejection. We provide a solution for this gap with ZK-Eval, studying code generation bottlenecks and evaluating end-to-end tasks.
Software Engineering for ZKP
The software engineering community has recently begun to explore tools and methodologies for ZKP development, with significant emphasis on security. Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024a), QED2 (Pailoor et al., 2023), CCC-Check (Kolozyan et al., 2025), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools (Xiao et al., 2025; Takahashi et al., 2025; Chaliasos et al., 2025) use fuzzing to expose security issues. These tools are invaluable for post-hoc analysis and verification, and can be integrated into our system for further security assurance. Our work complements these efforts by focusing on the generation phase, leveraging LLMs’ capabilities to reduce engineering burden.
7. Discussion and Limitations
ZKP Program Efficiency.
Program efficiency for ZKPs is measured by the size of arithmetic circuits, which directly determines prover cost and overall performance. To reduce circuit size, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce cost through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019). On the programmer side, several techniques further improve efficiency, such as adopting ZK-friendly hash functions (Grassi et al., 2021), gate packing, and aggregating bits into single field elements. For LLM-based code generation, however, producing efficient circuits is more challenging than generating correct ones. As future work, we plan to survey optimization techniques and guide LLMs on this.
LLM Training and Fine-tuning.
Recent studies show that LLMs struggle with DSLs and other low-resource languages due to the lack of high-quality training data (Joel et al., 2024), a challenge that is particularly acute in the context of ZKPs (Sheybani et al., 2025). This scarcity also limits the feasibility of fine-tuning approaches. Our work takes a step toward mitigating this limitation by systematizing the generation of program examples in ZK DSLs, which can support targeted training and domain adaptation of LLMs. These examples can serve as a complementary resource for future research in ZK code generation.
Real-World Representativeness.
We examine the representativeness and scope of our evaluation along two complementary dimensions. First, from a structural standpoint, the programs generated in § 5.4 exhibit an average cyclomatic complexity of 22.81, which exceeds the average observed in production-grade open-source ZK repositories (16.8 for Noir and 17.5 for Circom). This indicates that our benchmarks not only mirror but often surpass the structural depth of real-world ZK logic. Second, while our tasks are designed to isolate the core challenge of constraint specification, our evaluation is not “API-blind.” We recognize that LLM capabilities in standard and cryptographic API usage are already well established in prior work (Qin et al., 2023; Zhuo et al., 2024; Wang et al., 2025b; Zhu et al., 2025). Consequently, our test suite focuses on verifying that these APIs are integrated correctly within the circuit logic, covering a considerable number of real-world in-circuit invocations such as cryptographic primitives and data-processing utilities. This dual focus ensures that ZK-Coder is evaluated against the rigorous logical demands of production-level ZK development.
Threats to Validity.
We identify these potential threats. ➀ Our language-knowledge study uses multiple-choice QA items that were initially generated with LLM assistance. To mitigate bias and errors, we performed iterative human validation and cross-checking; while some residual noise may remain, we expect it to have limited impact on the reported trends. ➁ The ZK-Eval assessment uses hand-crafted test cases that exercise both accepting and rejecting behaviors of generated programs and primitive usages. Although tests cannot prove correctness, they include representative corner cases and explicit checks for soundness and completeness, which substantially reduce the risk of false positives. The high mutation detection rate reported in § 3.1.3, achieved through oracle-guided mutation testing, further supports this claim by demonstrating strong sensitivity to subtle constraint specification errors. ➂ The reliability of our automated evaluation depends on test-case coverage. We mitigate this through extensive cross-validation and edge-case injection, achieving a high mutation score. While this is not a formal proof of logical equivalence, it provides strong empirical assurance of correctness. ➃ The choice of prompting strategy may influence LLM performance. We employ standard zero-shot and few-shot prompting techniques without extensive prompt engineering or fine-tuning. To ensure experimental integrity, we verified that no test cases or ground-truth solutions were included in the prompts, preventing information leakage that could artificially inflate success rates. ➄ Security is critical for ZKPs. While integration with auditing tools (Pailoor et al., 2023; Kolozyan et al., 2025) could further enhance assurance, code generation from natural language specifications may face inherent ambiguity and underspecification. Users must exercise caution and validate generated code before deployment in security-critical application.
8. Conclusion
We presented ZK-Eval, the first benchmark for assessing LLM capability in ZK program generation, revealing that models struggle with composing algebraic primitives for verification constraints. To address this, we introduce ZK-Coder, which guides LLMs from constraint sketches to executable code through sketch-guided retrieval and iterative refinement, significantly improving code generation for both Circom and Noir.